# On some aspects of local thermodynamical formalism

### v2

@@ 4 Discussion on local entropy
Let $f:X\to X$ be a continuous map on a compact space. A set $E\subset X$ is an entropy set if at least two of these points in $E$ are distinct, and if for each open cover ${\mathcal{U}}=\{U_{1},\dots,U_{n}\}$ of $X$ with the property that for each $i$ there is $x\in E$ such that $x\notin\overline{U_{i}}$, the topological entropy $h_{\text{\tiny top}}(T,{\mathcal{U}})$ (in the sense of Adler-Konheim-McAndrew [2]) of the cover is positive. An entropy set of $n$ points is an entropy $n$-tuple, so an entropy pair is an entropy $2$-tuple.
Dou et al. [16] characterize maximal entropy sets, and prove that $E$ is an entropy set if and only if each $n$-tuple in $E$ with at least two distinct point is an entropy $n$-tuple.
The survey paper of Glasner & Ye [21] is largely on the relation between the topological and ergodic approach to dynamical systems from the global point of view. In particular, they discuss topological and measure entropy pairs and $n$-tuples and they present the relation between the topological entropy of open covers and the measure-theoretic entropy of finer partitions. The second part of the survey is on local properties of dynamical systems. They introduce $n$-tuples for an invariant measure, and establish their relation to topological entropy $n$-tuples. Glasner & Ye [21] discuss in detail the notion of entropy set, in order to find where the entropy is concentrated. In particular, it was shown that any topological dynamical system with positive entropy admits an entropy set with infinitely many points. For any topological dynamical system, there exists compact countable subset such that its Bowen entropy is equal to the entropy of the original system.
@@
In 1992, Thieullen [36] introduced a local form of entropy which he called $\alpha$-entropy. In his definition, the $\varepsilon$ in the Bowen $(n,\varepsilon)$-balls is replaced by an $n$-dependent quantity $e^{-\alpha n}$, but unlike our translocal entropy, he still uses Bowen balls. Ben Ovaria & Rodriguez-Hertz [4] introduced neutralized (local) entropy as
${\mathcal{E}}_{\mu}(x):=\lim_{\alpha\to 0}\limsup_{n\to\infty}-\frac{1}{n}\log\mu(B_{n}(x,e^{-\alpha n}));$ (11)
this is $\alpha$-entropy where eventually the limit $\alpha\to 0$ is taken. Ben Ovadia & Rodriguez-Hertz study the role it can play in hyperbolic dynamics and compare it with the Brin-Katok local entropy, see (6). They show that the neutralized local entropy for $C^{1+\beta}$ diffeomorphism $f:M\to M$ of a compact closed manifold $M$, calculated with respect to an $f$-invariant probability measure $\mu$, coincides with Brin-Katok local entropy almost everywhere.
In a recent paper, Yang, Chen & Zhou [38], extending (7), obtained a variational principle for neutralized entropy of a continuous map $f:X\to X$ on a compact metric space. For a non-empty compact $K\subset X$ they prove the equality
@@
Acknowledgments The first author was partially supported by the inner Łódź University Grant 11/IDUB/DOS/2021. This research was finalized during the visit of the second author to the Jagiellonian University funded by the program Excellence Initiative – Research University at the Jagiellonian University in Kraków.

### v3

@@ 4.2 Entropy pairs and tuples
Let $f:X\to X$ be a continuous map on a compact space $X$. A set $E\subset X$ is an entropy set if $E\geq 2$, and if for each open cover ${\mathcal{U}}=\{U_{1},\dots,U_{n}\}$ of $X$ with the property that for each $i$ there is $x\in E\setminus\overline{U_{i}}$, the topological entropy $h_{\text{\tiny top}}(f,{\mathcal{U}})$ (in the sense of Adler-Konheim-McAndrew [2]) of the cover is positive. An entropy set of $n$ points is an entropy $n$-tuple, so an entropy pair is an entropy $2$-tuple.
Dou, Ye & Zhang [19] characterize maximal entropy sets, and prove that $E$ is an entropy set if and only if each $n$-tuple in $E$ with at least two distinct point is an entropy $n$-tuple.
The survey paper of Glasner & Ye [25] is largely on the relation between the topological and ergodic approach to dynamical systems from the global point of view. In particular, they discuss topological and measure entropy pairs and $n$-tuples and present the relation between the topological entropy of open covers and the measure-theoretic entropy of finer partitions. The second part of the survey is on local properties of dynamical systems. They introduce $n$-tuples for an invariant measure, and establish their relation to topological entropy $n$-tuples. Glasner & Ye [25] discuss in detail the notion of entropy set, in order to determine where the entropy is concentrated, including the result of Dou, Ye and Zhang [19, Corollary 3.3], namely that any topological dynamical system with positive entropy admits an entropy set with infinitely many points. Moreover, for any topological dynamical system, there exists a compact countable subset such that its $(n,\varepsilon)$-separated set entropy is equal to the entropy of the whole system, see [49].
@@ 4.3 Neutralized entropy
In 1992, Thieullen [44] introduced a local form of entropy which he called $\alpha$-entropy and which is close to our approach of translocal entropy. In his definition, the $\varepsilon$ in the Bowen $(n,\varepsilon)$-balls is replaced by an $n$-dependent quantity $e^{-\alpha n}$, but unlike our translocal entropy, he still uses Bowen balls. Ben Ovadia & Rodriguez-Hertz [4] introduced neutralized (local) entropy as
${\mathcal{E}}_{\mu}(x):=\lim_{\alpha\to 0}\limsup_{n\to\infty}-\frac{1}{n}\log\mu(B_{n}(x,e^{-\alpha n})).$ (10)
This is the $\alpha$-entropy where eventually the limit $\alpha\to 0$ is taken. Ben Ovadia & Rodriguez-Hertz study the role it can play in hyperbolic dynamics and compare it with the Brin-Katok local entropy, see (7).
They show that the neutralized local entropy for $C^{1+\beta}$ diffeomorphism $f:M\to M$ of a compact closed manifold $M$, calculated with respect to an $f$-invariant probability measure $\mu$, coincides with Brin-Katok local entropy almost everywhere.
In a recent paper [48], Yang, Chen & Zhou, extending the variational principle (8) of Feng & Huang [21], obtained a variational principle for neutralized entropy of a continuous map $f:X\to X$ on a compact metric space. For a non-empty compact set $K\subset X$, they prove the equality
@@
[…]
Although neutralized entropy ${\mathcal{E}}_{\mu}(x)$ uses similar components as translocal entropy $h_{\omega}(x)$, they are hard to compare because the former depends on the measure $\mu$ and the latter on the parameter $\omega$.
Acknowledgments The first author was partially supported by the inner Łódź University Grant 11/IDUB/DOS/2021. This research was finalized during the visit of the second author to the Jagiellonian University funded by the program Excellence Initiative – Research University at the Jagiellonian University in Kraków. Both authors are very grateful for the careful reading and the many due corrections by the anonymous referees.